Can someone please explain me how to derive this answer.. please.. thanks..![Image]()
You do not have the required permissions to view the files attached to this post.
Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.
See classesLearn live with an expert or move at your own pace. Every option includes the complete TTP study system.
Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.
with Josh Braslow, 20+ years in test prep
Self-paced EA prep. Study on your schedule.
with
Complete access from day one. Study on your schedule.
Thanks for answereing, but the correct answer is 3/10 not 1/2. thanks again.. someone else please help me..freedsl wrote:they are asking for bread only buyers which equals neiter( C or M) buyers which equals 1-probability of M&C ONLY buyers.
C only = 50-20 =30
M only = 40-20 =20
then P (M&C) = 20+30/100 or 1/2
1-1/2 =1/2 probability of M only.
Kaching!jscoligan wrote:Cake+muffin+neither-both=Real total
50+40+neither-20=100
neither=30
30/100=> 3/10
